OpenStudy (anonymous):

If 1.6 moles of an ideal gas are at a temperature of 25°C and a pressure of 0.91 atm Hg, what volume would the gas occupy, in liters? 0.28 L 3.0 L 36 L 43 L @nechirwan

OpenStudy (nechirwan):

Just use the ideal gas law: PV = nRT. P = pressure in atm V = volume in liters n = moles of gas R = gas constant = 0.0821 L atm / K mole T = Kelvin temperature = 25°C + 273 = 298 K PV = nRT V = nRT / P = (1.6)(0.0821)(298) / (0.91) = 43 L